95/100+8/1000=?
pls answer me step by steps

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

95/100+8/1000 = 950/1000 + 8/1000
= 958/1000
= 479/500

Answer by stanbon(75887)

95/100+8/1000=?
-------
least common denominator: 1000
-----
(95/100)(10/10) + (8/1000)
-----
= (950/1000)+(8/1000)
--------
= (950+8)/1000
--------
= 958/1000
====================
Cheers,
Stan H.
====================